# Structure of Classifier Boundaries: Case Study for a Naïve Bayes Classifier

### v2

@@ 4 Results
All analyses reported here were performed using the R Language and Environment for Statistical Computing (R Core Team,, 2020). Data and R code are available at https://doi.org/10.7910/DVN/FXOFLI.
@@ 4.1 DNA Reads
Table 2 contains cross-tabulations of Boundary Status for the 5869 reads as a function of (1) read source, (2) the classifier decision and (3) correctness of the classifier decision. Before discussing details, we observe that 1791 of the 5869 reads, or 30.52%, lie on the boundary $\mathcal{B}$, the first concrete evidence that $\mathcal{B}$ is not “thin.” Of these 1791 boundary reads, 190 (3.24%) have Boundary Status equal to 2.
Looking in more detail, Figure 1 shows the neighbor distributions over $\mathcal{O}$ for all the reads. In it, three-dimensional probabilities (barycentric coordinates) are converted to Cartesian coordinates, as points in an equilateral triangle. Pure Adeno, in the sense that $\mathrm{Prob}(\mbox{Adeno})=1$, is the top vertex, pure COVID is the lower left vertex, and pure SARS is the lower right vertex. Because we know the sources of the reads, we create separate displays for each source. The white diamond in each scatterplot is the centroid of the probabilities it contains, and all three centroids lie close to the source. A read on an edge has neighbors of two kinds, one of which is that of the read itself. (In theory, a read could have $\mbox{BS}(R)=2$ without neighbors of its same type, but this bizarre case of a read classified differently from all of its neighbors does not occur in our dataset. The minimum number of concordant neighbors is 11.) A read in the interior of the triangle has neighbors with all three decisions.
@@
The middle table in Table 2 provides a complementary view of Boundary Status as a function of classifier decision, so that its row sums match the column sums in in Table 1. Here also, the $\chi^{2}$ test is massively significant. SARS remains the outlier: relatively more reads classified as SARS lie on the boundary than for Adeno or COVID.
The bottom table in Table 2 shows Boundary Status as a function of correctness of the classifier decision, and begins to get at the heart of the matter. Of incorrectly decided reads, 64.66% lie on the boundary, while only 22.81% of correctly decided reads lie on the boundary! The classifier is struggling in the vicinity of $\mathcal{B}$, and one rightly should be less confident of such decisions. The $\chi^{2}$ statistic for this table is 726.65, more than five times that for source (topmost table), and three times that for decision (middle table). Moreover, the Kolmogorov-Smirnov test for the two empirical cumulative distribution functions (ECDFs) is massively significant: Boundary Status is higher for incorrect decisions than for correct ones.
Figure 1: Neighbor distributions for the 5869 reads, by source.
@@
In parallel to Table 2, Figures 2 shows the distribution of Neighbor Similarity as a function of read source, read decision and decision correctness. In particular, these ECDFs make clear how many reads are not on the boundary, for which $\mbox{NS}(R)<1$. The qualitative messages mirror those in Table 2.
Figure 2: ECDFs of Neighbor Similarity. Left: by read source. Center: by classifier decision. Right: by decision correctness
Neighbor Similarity is, obviously, more burdensome computationally than simply running the classifier. However, the computation is infinitely parallelizable: classifier evaluation for one neighbor is completely independent of that for other neighbors. An alternative strategy is to sample neighbors, which generates estimates for Neighbor Similarity. Figure 3 provides initial evidence that sampling can be effective. It is based on the entire 5869-element read dataset. The $x$-axis is the number of reads sampled; the order was randomized for each read. The $y$-axis is more complicated. For each read $R$ and sample size $k$ we calculated a partial Neighbor Similarity $\mbox{NS}(R,k)$ using (6). Then for each $k$, we fitted a linear model with $\mbox{NS}(\cdot,404)=\mbox{NS}$ as response and $\mbox{NS}(\cdot,1),\dots,\mbox{NS}(\cdot,k)$ as predictors. Finally, the $y$-axis contains values of relative root mean squared error (RRMSE) [^8] The square root of the mean squared error for the linear model divided by the mean response. for each model. From the figure, an RRMSE of 5% requires only 20 sampled neighbors, and there is minimal benefit from sample sizes exceeding 80. How these values scale with problem size is not clear.
Figure 3: Relative root mean squared error in estimating NS from samples of neighbors.
4.2 Relationship to Inherent Measures of Uncertainty
We next investigate the relationships between Neighbor Similarity and the two inherent measures for $C$: $\mbox{MP}(\cdot)$—the maximum of the three posterior probabilities, and $\mbox{PE}(\cdot)$—the entropy of the posterior distribution. Always, $\mbox{MP}(R)\geq 1/3$, and the closer $\mbox{MP}(R)$ is to 1, the more certain $C$ is of the decision for $R$. Figure 4 depicts the relationship between MP and NS as a function of classifier decision. Because one ultimate goal is to employ NS in contexts where neither MP nor a model-derived analog is available, it would not have been appropriate to use read source there rather classifier decision. The correlation between MP and NS is $0.8488$; the disaggregated correlations are 0.8784 for Adeno, 0.8502 for COVID and 0.8400 for SARS.
@@ 4.2 Relationship to Inherent Measures of Uncertainty
As the notation suggests, the process is equivalent to fitting three separate models to datasets constructed by filtering on $C(R)$. The adjusted coefficient of determination for the model is 0.8341, indicating good fit. Figure 5 shows actual and predicted values of MP by classifier decision. Much of the imperfect fit is simple inability to predict the continuous variable MP within sets of reads having the same value of NS (especially those with $\mbox{NS}=1$), which the figure makes perfectly clear.
Figure 4: Scatterplots of MP versus NS, by classifier decision. Left: decision = Adeno, Center: decision = COVID. Right: decision = SARS.
@@
The second inherent measure of uncertainty for $C$ is posterior entropy PE. The correlations between NS and PE are -0.7167 (aggregated), -0.7437 (Adeno), -0.7280 (COVID) , and -0.6964 (SARS). Analogously to Figure 4, Figure 7 contains scatterplots of PE versus NS by classifier decision. Clearly, PE is predicted less well by NS and classifier decision than is MP. For a quadratic regression model analogous to that in (8), the adjusted $R^{2}$ is only 0.6928. The corresponding partition model, shown in Figure 8, identifies which variables and interactions are most relevant. Again, the mean squared error is slightly lower than for the quadratic regression model.
@@ 4.3 Exploring the Boundary
The first and simplest method for locating $\mathcal{B}$ has been mentioned already in Section 3.1: if $C(R)\neq C(R^{\prime})$, then any path from $R$ to $R^{\prime}$ contains at least two elements of $\mathcal{B}$. Within the reads dataset, there are 11,480,223 such pairs. If we eliminate duplicates and for each pair consider only the Hamming path that in left-to-right order replaces each nucleotide in $R$ differing from the corresponding nucleotide in $R^{\prime}$ by the latter, then we have a mechanism for locating boundary pairs. Based on a random sample of 25 origin reads, but considering all other reads with different decisions, 1,242,162 boundary points were identified. There is significant variability, however, as shown in Table 3. Moreover, the table shows that paths originating from reads classified as COVID differ significantly from those originating at reads classified as Adeno or SARS.
@@
Figure 9: Numbers of boundary points on Hamming paths, by classifier decision.
Of course, whenever $C(R)\neq C(R^{\prime})$, there are multiple Hamming paths connecting $R$ and $R^{\prime}$. Initial experiments have shown that a strategy of using multiple Hamming paths is not materially more effective than using one Hamming path in terms of unique boundary pairs identified relative to computational effort. The reason, we hypothesize, is that multiple paths from $R$ to $R^{\prime}$ often cross $\mathcal{B}$ at the same places.
A second strategy, motivated by Markov chain Monte Carlo (MCMC) methods used to investigate large spaces of contingency tables (Diaconis and Sturmfels,, 1998), is to perform random walks originating from sequences $x\in\mathcal{I}$, moving to a neighboring sequence at each step, and checking whether $\mathcal{B}$ has been crossed, i.e., whether the classifier decision has changed. Were $\mathcal{B}$ thin, this strategy of wandering randomly would be horribly ineffective. But, because $\mathcal{B}$ is not thin, it works reasonably well. To illustrate, we chose a random sample of 100 reads for each decision, and ran random walks of length 2000 steps originating at each. Keeping in mind that each boundary pair comprises two boundary points, 35,352 boundary points were found, of 600,300 sequences visited.
@@
Note that in our implementation both Hamming paths and random walks may pass through boundary points without identifying them, which can happen because, for computational efficiency, each point in the path is compared only to its predecessor rather than all of its neighbors.
Finally, we present random walk-like methods for “crawling the boundary” once it has been reached. We simulated 100 random walks starting at boundary points, but constrained to move only to adjacent and not previously visited boundary points. Each crawl was allowed to continue for 250 steps, or until it was not possible to move to another and not previously visited boundary point. The fact that such points exist further confirms the geometric complexity of the boundary: it is ”hairy” in the sense that the points from which no continuation is possible are tips of hairs.
Crawling the boundary, while insightful, is not efficient. Of the 100 random walks, 64 did not terminate within 250 steps, indicating that the boundary $\mathcal{B}$ is significantly contiguous. The lengths of crawls that did terminate range from 5 to 242, with no discernible structure to their distribution. The 36 crawls that did terminate are the hairs mentioned in the preceding paragraph—points at which continuing on the boundary is possible only by backtracking. While in the minority, they are nevertheless not rare. Beyond this, the boundary is convoluted: all three decisions are present in 58 of the 100 crawls. Put differently, if one were to think of three boundaries—an Adeno–COVID boundary, an Adeno–SARS boundary, and a COVID–SARS boundary—they cross each other frequently, as evidenced by the 190 reads with neighbors of all three types. In all other crawls, the decisions must and do alternate, indicating that some segments of the boundary lie between two decisions.
The inefficiency of crawling the boundary, which is apparent in Table 5, is partly an artifact of our computational implementation. At each step, we classify all neighbors before determining if it is possible to proceed. Computation would be reduced by examining neighbors sequentially, and proceeding once a neighbor with a differing classification is identified.
@@ 4.4 Do Boundary Points Differ from Other Reads?
Our final results address a natural question: do the 1791 boundary reads in our dataset differ from the 4078 other reads? Figure 10 provides initial insight. It contains scatterplots of the four components of a multi-dimensional scaling (MDS) (Kruskal and Wish,, 1978) of the triplet distributions, colored by Boundary Status. The most striking feature is the concentration of values of Component 1 for reads $R$ with $\mbox{BS}(R)=2$, which is reinforced by the upper left-hand panel in Figure 11. This latter figure contains plots of the ECDFs of each of the four components, disaggregated by Boundary Status; the colors match those in Figure 10. In fact because of overplotting in Figure 10, there is additional structure that is less visible in it than in Figure 11. Notably, there are many instance where one ECDF lies entirely below another, including Component 1 ($\mbox{BS}=2$) below Component 1 ($\mbox{BS}=1$); effectively, Component 3 ($\mbox{BS}=0$) below Component 3 ($\mbox{BS}=1$) below Component 3 ($\mbox{BS}=2$); Component 4 ($\mbox{BS}=0$) below Component 4 ($\mbox{BS}=1$); and Component 4 ($\mbox{BS}=2$) below Component 4 ($\mbox{BS}=1$). Almost uniformly the differences among the ECDFs, as evaluated by Kolmogorov-Smirnov tests, are statistically significant.
@@
Many analysts would still be uncomfortable that this model is overfit, so in Table 7, we show all possible pruned versions of the model, as generated by the R package rpart. At the top is the “default model” that predicts all reads to have Boundary Status 0. At the bottom is the unpruned, and perfect, model. Halving the incorrect classification rate from 30% to 15%, for instance, requires only 182 terminal nodes. Their sizes range from 3 to 445, with a mean size of 32 and median of 14. To demonstrate the scientific insights that follow from this approach, Table 10, in Appendix B, shows the importance of each of the 64 predictors to the splits that comprise the model, as percentages that sum to 100%. In order, the five most important predictors are TTT (which codes for the amino acid phenylalanine), TAT (tyrosine), GTT (valine), GCG (alanine), and AAG (lysine).
Finally, regarding validation, if the 705-node pruned model is applied to the 6000-element ReadsNew dataset discussed in Section 5.2.1, it correctly predicts Boundary Status in 4464 cases, or 74.2%, which is respectable performance given the limited training data.
@@
Table 7: For all possible pruned partition models, the number of terminal nodes and correct classification rate. The principal pruned version is indicated by ***, and the “halve the error rate” model by $+$

### v3

@@ 4 Results
All analyses reported here were performed using the R Language and Environment for Statistical Computing (R Core Team,, 2020).
@@ 4.1 DNA Reads
Table 2 contains cross-tabulations of Boundary Status for the 5869 reads as a function of (1) read source, (2) the classifier decision and (3) correctness of the classifier decision. We observe that 1791 of the 5869 reads, or 30.52%, lie on the boundary $\mathcal{B}$, the first concrete evidence that $\mathcal{B}$ is not “thin.” Of these 1791 boundary reads, 190 (3.24%) have Boundary Status equal to 2.
Looking in more detail, Figure 1 shows the neighbor distributions over $\mathcal{O}$ for all the reads. In it, three-dimensional probabilities (barycentric coordinates) are converted to Cartesian coordinates, as points in an equilateral triangle. Pure Adeno, in the sense that $\mathrm{Prob}(\mbox{Adeno})=1$, is the top vertex, pure COVID is the lower left vertex, and pure SARS is the lower right vertex. Because we know the sources of the reads, we create separate displays for each source. The white diamond in each scatterplot is the centroid of the probabilities it contains, and all three centroids lie close to the source. A read on an edge has neighbors of two kinds, one of which is that of the read itself. (In theory, a read could have $\mbox{BS}(R)=2$ without neighbors of its same type, but, as noted, this case of a read classified differently from all of its neighbors does not occur in our dataset. The minimum number of concordant neighbors is 11.) A read in the interior of the triangle has neighbors with all three decisions.
@@
The middle table in Table 2 provides a complementary view of Boundary Status as a function of classifier decision, so that its row sums match the column sums in Table 1. Here also, the $\chi^{2}$ test is massively significant, with $p$-value less than $10^{-16}$. SARS remains the outlier: relatively more reads classified as SARS lie on the boundary than for Adeno or COVID. We do not have a biological hypothesis for this.
The bottom table in Table 2 shows Boundary Status as a function of correctness of the classifier decision, and begins to get at the heart of the matter. Of incorrectly decided reads, 64.66% lie on the boundary, while only 22.81% of correctly decided reads lie on the boundary! The classifier is struggling in the vicinity of $\mathcal{B}$, and one rightly should be less confident of such decisions. The $\chi^{2}$ statistic for this table is 756.86, more than five times that for source (topmost table), and three times that for decision (middle table). Moreover, the Kolmogorov-Smirnov test for the two empirical cumulative distribution functions (ECDFs) is massively significant, with a p-value less than $10^{-16}$: Boundary Status is higher for incorrect decisions than for correct ones.
Figure 1: Neighbor distributions for the 5869 reads, by source. The geometry is explained in the text.
@@
In parallel to Table 2, Figure 2 shows the distribution of Neighbor Similarity as a function of read source, read decision and decision correctness. In particular, these ECDFs make clear how many reads are not on the boundary, for which $\mbox{NS}(R)<1$. The qualitative messages mirror those in Table 2.
Figure 2: ECDFs of Neighbor Similarity. Left: by read source. Center: by classifier decision. Right: by decision correctness. In each panel, the $x$-axis is Neighbor Similarity and the $y$-axis is cumulative probability.
Neighbor Similarity is more burdensome computationally than simply running the classifier. However, the computation is infinitely parallelizable: classifier evaluation for one neighbor is completely independent of that for other neighbors. An alternative strategy is to sample neighbors, thereby generating estimates for Neighbor Similarity. Figure 3 provides initial evidence that sampling can be effective. It is based on the entire 5869-element read dataset. The $x$-axis is the number of reads sampled; the order was randomized for each read. The $y$-axis is more complicated. For each read $R$ and sample size $k$ we calculated a partial Neighbor Similarity $\mbox{NS}(R,k)$ using (7). Then for each $k$, we fitted a linear model with $\mbox{NS}(\cdot,404)=\mbox{NS}$ as response and $\mbox{NS}(\cdot,1),\dots,\mbox{NS}(\cdot,k)$ as predictors. Finally, the $y$-axis contains values of relative root mean squared error (RRMSE—the square root of the mean squared error for the linear model divided by the mean response) for each model. From the figure, an RRMSE of 5% requires only 20 sampled neighbors, and there is minimal benefit from sample sizes exceeding 80. How general these results are, and in particular how values scale with respect to problem size, is not clear.
Figure 3: Relative root mean squared error (RRMSE) in estimating NS from samples of neighbors. The $x$-axis is the number of samples and the $y$-axis is RRMSE.
4.2 Relationship of Neighbor Similarity to Inherent Measures of Uncertainty
We next investigate the relationships between Neighbor Similarity and the two inherent measures for $C$: $\mbox{MP}(\cdot)$—the maximum of the three posterior probabilities, and $\mbox{PE}(\cdot)$—the entropy of the posterior distribution.
Always, $\mbox{MP}(R)\geq 1/3$, and the closer $\mbox{MP}(R)$ is to 1, the more certain $C$ is of the decision for $R$. Figure 4 depicts the relationship between MP and NS as a function of classifier decision. Because one ultimate goal is to employ NS in contexts where neither MP nor a model-derived analog is available, it would not have been appropriate to use read source there, rather than classifier decision. The correlation between MP and NS is $0.8488$; the Spearman correlation is 0.7945. Disaggregated correlations are 0.8784 for Adeno, 0.8502 for COVID and 0.8400 for SARS, and corresponding Spearman values are 0.8784, 0.8502, and 0.8310.
@@ 4.2 Relationship of Neighbor Similarity to Inherent Measures of Uncertainty
As the notation suggests, the process is equivalent to fitting three separate models to datasets constructed by filtering on $C(R)$. The adjusted coefficient of determination for the model is 0.8341, indicating good fit. Figure 5 shows actual and predicted values of MP by classifier decision. Much of the imperfect fit results from inability to predict the continuous variable MP within sets of reads having the same value of NS (especially those with $\mbox{NS}=1$), which the figure makes perfectly clear.
Figure 4: Scatterplots of MP versus NS, by classifier decision. Left: decision = Adeno, Center: decision = COVID. Right: decision = SARS. In each panel, the $x$-axis is Neighbor Similarity and the $y$-axis is MP.
@@
The second inherent measure of uncertainty for $C$ is posterior entropy PE. The correlations between NS and PE are -0.7167 (aggregated), -0.7437 (Adeno), -0.7280 (COVID) , and -0.6964 (SARS). Analogously to Figure 4, Figure 7 contains scatterplots of PE versus NS by classifier decision. Clearly, PE is predicted less well by NS and classifier decision than is MP. For a quadratic regression model analogous to that in (8), the adjusted $R^{2}$ is 0.6928. The corresponding partition model, shown in Figure 8, identifies which variables and interactions are most relevant. Again, the mean squared error is slightly lower than for the quadratic regression model.
@@ 4.3 Exploring the Boundary
The first and simplest method for locating $\mathcal{B}$ has been mentioned in Section 3.1: if $C(R)\neq C(R^{\prime})$, then any path from $R$ to $R^{\prime}$ contains at least two (neighboring) elements of $\mathcal{B}$. Within the reads dataset, there are 11,480,223 such pairs. If we eliminate duplicates and for each pair consider only the Hamming path that in left-to-right order replaces each nucleotide in $R$ differing from the corresponding nucleotide in $R^{\prime}$ by the latter, then we have a mechanism for locating boundary pairs. (As noted below, this particular choice of Hamming path is inconsequential.) Based on a random sample of 25 origin reads, but considering all other reads with different decisions, 1,242,162 boundary points were identified. There is significant variability, however, as shown in Table 3. Moreover, the table shows that paths originating from reads classified as COVID differ from those originating at reads classified as Adeno or SARS.
@@
Of course, whenever $C(R)\neq C(R^{\prime})$, there are multiple Hamming paths connecting $R$ and $R^{\prime}$. Initial experiments have shown that a strategy of using multiple Hamming paths is not materially more effective than using one Hamming path in terms of unique boundary pairs identified relative to computational effort. The reason, we believe, is that multiple paths from $R$ to $R^{\prime}$ often cross $\mathcal{B}$ at the same places.
A second strategy, motivated by Markov chain Monte Carlo (MCMC) methods used to investigate large spaces of contingency tables (Diaconis and Sturmfels,, 1998), is to perform random walks originating from sequences $x\in\mathcal{I}$, moving to a neighboring sequence at each step, and checking whether $\mathcal{B}$ has been crossed, i.e., whether the classifier decision has changed. Were $\mathcal{B}$ ‘thin,” this strategy of wandering randomly would be horribly ineffective. But, because $\mathcal{B}$ is not thin, it works reasonably well. To illustrate, we chose a random sample of 100 reads for each decision, and ran random walks of length 2000 steps originating at each. Keeping in mind that each boundary pair comprises two boundary points, 35,352 boundary points were found, of 600,300 sequences visited.
@@
Note that in our implementation both Hamming paths and random walks may pass through boundary points without identifying them, which can happen because (for computational efficiency) each point in the path is compared only to its predecessor rather than all of its neighbors.
Finally, we present random walk-like methods for “crawling the boundary” once it has been reached. We simulated 100 random walks starting at boundary points, but constrained to move only to adjacent and not previously visited boundary points. Each crawl was allowed to continue for 250 steps, or until it was not possible to move to another and not previously visited boundary point. The fact that such points exist further illuminates the geometric complexity of the boundary: it is ”hairy” in the sense that the points from which no continuation is possible are tips of hairs.
While insightful, crawling the boundary is not efficient. Of the 100 random walks, 64 did not terminate within 250 steps, indicating that the boundary $\mathcal{B}$ is significantly contiguous. (See also §4.5.) The lengths of crawls that did terminate range from 5 to 242, with no discernible structure to their distribution. The 36 crawls that did terminate are the hairs mentioned in the preceding paragraph—points at which continuing on the boundary is possible only by backtracking. While in the minority, they are nevertheless not rare.
Beyond this, the boundary is convoluted: all three decisions are present in 58 of the 100 crawls. Put differently, if one were to think of three boundaries—an Adeno–COVID boundary, an Adeno–SARS boundary, and a COVID–SARS boundary—they cross each other frequently, as evidenced by the 190 reads with neighbors of all three types. In all other crawls, the decisions must and do alternate, indicating that some segments of the boundary lie between two decisions.
The inefficiency of crawling the boundary, which is apparent in Table 5, is partly an artifact of our computational implementation. At each step, we classify all neighbors before determining if it is possible to proceed. For related discussion, see Section 4.5.
@@ 4.4 Do Boundary Points Differ from Other Reads?
Here we address a natural question: do the 1791 boundary reads in our dataset differ from the 4078 other reads? Figure 9 provides initial insight. It contains scatterplots of the four components of a multi-dimensional scaling (MDS) (Kruskal and Wish,, 1978) of the triplet distributions, colored by Boundary Status. The most striking feature is the concentration of values of Component 1 for reads $R$ with $\mbox{BS}(R)=2$, which is reinforced by the upper left-hand panel in Figure 10. This latter figure contains plots of the ECDFs of each of the four components, disaggregated by Boundary Status; the colors match those in Figure 9. In fact because of overplotting in Figure 9, there is additional structure that is less visible in it than in Figure 10. Notably, there are many instances where one ECDF lies entirely below another, including Component 1 ($\mbox{BS}=2$) below Component 1 ($\mbox{BS}=1$); effectively, Component 3 ($\mbox{BS}=0$) below Component 3 ($\mbox{BS}=1$) below Component 3 ($\mbox{BS}=2$); Component 4 ($\mbox{BS}=0$) below Component 4 ($\mbox{BS}=1$); and Component 4 ($\mbox{BS}=2$) below Component 4 ($\mbox{BS}=1$).
@@
Many analysts would still be uncomfortable that this model is overfit, so in Table 7, we show all possible pruned versions of the model, as generated by the R package rpart. At the top is the “default model” that predicts all reads to have Boundary Status 0. At the bottom is the unpruned, and perfect, model. Halving the incorrect classification rate from 30% to 15%, for instance, requires only 182 terminal nodes. Their sizes range from 3 to 445, with a mean size of 32 and median of 14. The first column in this table contains the complexity parameter, which quantifies the relationship between model complexity and model fit. More precisely, it is the improvement in model fit resulting from making additional splits. The table contains only values at which the structure of the optimal tree changes.
Finally, regarding validation, if the 705-node pruned model is applied to the 6000-element “Validation” dataset discussed in Section 4.6.1, it correctly predicts Boundary Status in 4464 cases, or 74.2%, which is respectable performance given the limited training data.
@@
Table 7: For all possible pruned partition models, the number of terminal nodes and correct classification rate. The principal pruned version is indicated by ***, and the “halve the error rate” model by +++
@@
4.5 Is the Boundary Connected?
It is natural to ask whether the boundary is connected, in the sense that for any two points on it, there is a path connecting them that consists entirely of boundary points. It is clear that a definitive answer is not feasible computationally. Our evidence to date is that, at least in the vicinity of the training data, the boundary is connected. Of multiple attempts to connect boundary points using a greedy algorithm, none failed.
The algorithm appears in Algorithm 1. Translated into English, to construct a path from $R_{o}\in\mathcal{B}$ to $R_{d}\in\mathcal{B}$ ($o$ is for origin, $d$ is for destination) that does not leave $\mathcal{B}$, simply change the classifier decision at each step, which ensures remaining within $\mathcal{B}$, and reduce the distance to $R_{d}$. If this is not possible, remain the same distance from $R_{d}$. And if this is not possible, stop and return an error. The latter was never necessary, among nearly 100 cases that varied with respect to read source, classifier decision, and correctness of the decisions for $R_{o}$ and $R_{d}$.
Algorithm 1 Algorithm for connecting boundary points without leaving the boundary.
1:Bayes classifier $C$ and boundary points $R_{o},R_{d}\in\mathcal{B}$
[…]
3:Path to be constructed: $BP=(R_{o})$
4:Current point on path: $CP=R_{o}$
5:Current distance to $R_{d}$: $CD=H(CP,R_{d})$, where $H$ = Hamming distance
6:while $CP\neq R_{d}$ do
7:  Let $N(CP)$ be the set of all neighbors of CPCP
8:  $X\leftarrow\left\{R\in\mathcal{N}(CP):R\notin BP,C(R)\neq C(CP)\right\}$
9:  $X_{1}\leftarrow\left\{R\in R:H(R^{*},R_{d})<H(CP,R_{d})\right\}$
10:  if $X_{1}\neq\emptyset$ then
11:   Select $R^{*}$ randomly from $X_{1}$
12:   $BP\leftarrow(BP,R^{*})$
13:   $CP\leftarrow R^{*}$
14:  else if $X\neq\emptyset$ then
15:   Select $R^{*}$ randomly from $X$
16:   $BP\leftarrow(BP,R^{*})$
17:   $CP\leftarrow R^{*}$
18:  else
19:   return Error
20:  end if
21:end while
22:return BPBP
4.6 Scientific Generalizability
There are, of course, questions of scientific generalizability: are the properties observed somehow specific to our read dataset, or descriptive of the classifier $C$ more generally? And do they apply to other classifiers? Do they extend to other forms of data?
4.6.1 Other Datasets
To address the first question, we ran the Bayes classifier on a number of additional datasets, which are described in Table 8. Results appear in Table 9. For completeness, the training dataset itself is in the first row. All genomes were downloaded from NCBI.
These alternative datasets fall into four classes:
Validation:
This dataset is effectively a clone of the training dataset, and has been used in Porter and Karr, (2024), Karr and Ruane, (2026), and elsewhere to to evaluate model performance, which should (and does) match that for the training data.
Random sequences,
to possibly reach regions of $\mathcal{I}$ in which “real” DNA sequences are sparse.
Local behavior,
in Distance2, examining near-neighbors of Validation.
Other genomes,
which include other viruses, bacteria and the human genome.
For each, there are both one dataset of the nature just described and a second constructed by randomly sampling Hamming paths between 400 randomly selected pairs, which represent a more localized sampling of $\mathcal{I}$. See also Section 4.3.
Most important, all of the alternative datasets can be regarded as methods for sampling randomly from $\mathcal{I}$.
Table 8: Short descriptions of alternative datasets.
Name Description Size
Training Used throughout 5869
Validation 2000 Mason simulator-generated length-101 reads each from Adeno, COVID and SARS genomes 6000
Random Random sequences of length 101 10000
Random2 Random sequences of length 100 25000
Distance2 Two randomly generated Hamming distance 2 neighbors of each elements of Validation 12000
3Viruses Mason simulator-generated reads from norovirus, human and H5N1 influenza virus 1713
Ecoli Length 101 Mason simnulator-generated reads from an Escherichia coli genome 12000
Pg Length 101 Mason simnulator-generated reads from an Porphyromonas gingivalis genome 25000
Hsc1 Length 101 Mason simnulator-generated reads from an Home sapiens chromosome 1 25000
TrainingPaths All elements of a random sample of Hamming paths between 400 randomly selected pairs of Training 30188
ValidationPaths All elements of a random sample of Hamming paths between 400 randomly selected pairs of Validation 30452
RandomPaths All elements of a random sample of Hamming paths between 400 randomly selected pairs of Random 30697
Random2Paths All elements of a random sample of Hamming paths between 400 randomly selected pairs of Random2 30590
Distance2Paths All elements of a random sample of Hamming paths between 400 randomly selected pairs of Distance2 30343
3VirusesPaths All elements of a random sample of Hamming paths between 400 randomly selected pairs of 3Viruses 30440
EcoliPaths All elements of a random sample of Hamming paths between 400 randomly selected pairs of Ecoli 30713
PgPaths All elements of a random sample of Hamming paths between 400 randomly selected pairs of Pg 30776
Hsc1Paths All elements of a random sample of Hamming paths between 400 randomly selected pairs of Hsc1 30497
Boundary Status distributions for each dataset in Table 8 appear in Table 9. The most striking feature is that only for the Ecoli datasets does less 20% of the dataset lie on the boundary, and then only slightly. We interpret this as further evidence supporting our claim that the boundary is large. Moreover, for some datasets, one-half of the points lie on the boundary.
The results for Distance2, and to a lesser extent 3Viruses, show an important insight that at first may seem paradoxical. The boundary is denser in the vicinity of the Training and Validation, and sparser for Random, Random2, Ecoli, Pg and Hsc1. The reason for this, as discussed in detail and from a slightly different perspective in Karr and Ruane, (2026), is that in regions distant from the training data the classifier simply degenerates. In this case, as demonstrated by the “Frac(Adeno)” column in Table 9, nearly everything that is distant from the training data is classified as Adeno. Hence, there is no paradox: near the training data, we want high discrimination, which means a larger boundary. Also, therefore, Boundary Status is usable as a diagnostic indicating when a classifier is being used on data that differ from what was “expected.” For the Bayes classifier, “what is expected” is incorporated in the prior and the likelihood functions (Karr et al.,, 2023). For other classifiers, the training data represent “what is expected.” Finally, of course, one must temper interpreting high Neigbhor Similarity as confidence. This works only when the analysis data are not too far from “expected.”
For full disclosure, we note that Hsc1 does not differ from the training data nearly as much as do Random, Random2, Ecoli and Pg. We are not able to explain this, nor to hypothesize a plausible explanation for it.
Finally, we note that the material here complements that in Section 4.1 in that they are alternative methods to explore the boundary.
Table 9: Boundary Status Distributions for Bayes classifier applied to alternative datasets. Columns Boundary Status are percentages. The column Frac(Adeno) is fraction of predictions that are Adenovirus.
BoundaryStatus Frac(Adeno)
Dataset 0 1 2
Training 0.695 0.273 0.032 0.329
Validation 0.687 0.282 0.03 0.324
Random 0.792 0.196 0.012 0.912
Random2 0.789 0.201 0.01 0.911
Distance2 0.672 0.293 0.034 0.334
3Viruses 0.48 0.489 0.031 0.472
Ecoli 0.877 0.107 0.017 0.928
Pg 0.76 0.218 0.023 0.856
Hsc1 0.554 0.381 0.064 0.456
TrainingPaths 0.575 0.352 0.073 0.348
ValidationPaths 0.594 0.346 0.06 0.384
RandomPaths 0.798 0.194 0.009 0.916
Random2Paths 0.792 0.196 0.012 0.91
Distance2Paths 0.589 0.336 0.075 0.4
3VirusesPaths 0.49 0.465 0.044 0.556
EcoliPaths 0.828 0.152 0.021 0.916
PgPaths 0.752 0.224 0.024 0.868
Hsc1Paths 0.514 0.406 0.08 0.487
4.6.2 Other Classifiers
Full results will be reported elsewhere (Karr and Ruane,, 2026), but initial experiments show similar behavior for other classifiers applied to the same dataset, including partition models, random forests and neural networks (Porter and Karr,, 2024). Table 10, which also appears there, shows that performance of four classifiers on the training data, with respect to correct classification rate, size of the boundary and mean Neighbor Similarity.
Table 10: Comparison of four classifiers. CCR = correct classification rate; $|\mathcal{B}|$ = size of boundary; $\overline{NS}$ = mean Neighbor Similarity.
Classifier CCR $|\mathcal{B}|$ $\overline{NS}$
Bayes .8155 1791 .9292
Partition model .8075 4484 .8393
Random forest 1.0000 2173 .9897
Neural net .7657 2154 .9022
4.6.3 Other Forms of Data
The only essential characteristics of our framework are the graph structure of $\mathcal{I}$, which requires discreteness, and discreteness of $\mathcal{O}$. To illustrate, $\mathcal{I}$ could be the space of all $1024\times 1024$ RGB images with 8-bit color depth. Therefore, no one should be surprised that there are images that look the same but are classified differently. Indeed, because there is a boundary, there are images that differ by one color level in one pixel that are classified differently. Of course in this setting, computational issues preclude anything remotely resembling the analyses here.